-- HarryRenshall - 16 Jan 2008

Week of 080114

Open Actions from last week:

Monday:

see the weekly phone conference in Indico

Tuesday:

EIS-on-Duty report: ALICE (PM) met with the Castor team this morning to discuss in which Castor release will be a new xrootd plugin that they need. More info tomorrow. Since RAL is staying with Castor 2.1.4 (this and 2.1.6 are supported for bug fixes) ALICE will send raw data to tape at RAL in February but not read it back. ATLAS (SC) software to handle space tokens is ready and will be used as soon as sites configure them. CMS (AS) nothing to report. LHCb (RS) Nick Brook will coordinate solving the Castor/rfio problem of LGCb at RAL and CNAF. There will be an LFC bulk data modification to the master LFC.

SMOD report: (M.C-S) FIO will upgrade all CERN LFC to 1.6.7 next week. The 1.6.8 version which runs in slc4 is not yet available though ATLAS plan to use it for bulk operations in February.

DBMOD report: (JDS) We now know that the physics databases client on worker nodes needs much of the applications area software stack. It also needs Oracle libraries and these are in fact freely available but without technical support. Migration to new hardware in the integration RAC will be done for CMS this week and ATLAS next week.

Monitoring / dashboard report: We are looking at how best to handle the critical services.

Release update: Nothing new to report.

Questions from sites:

AOB: D.Bonacorsi of CMS asked if Castor sites are free to choose if they run 2.1.4 or 2.1.6 ? (M.C-S) Yes, both are supported for bug fixes. Also can they choose gridftp1 or 2 ? (M.C-S) Castor 2.1.6 will require gridftp2. (JS) the approved middleware versions should be ready to push out to sites next week. (NT) the SL4 VO-box is ready to go into production - are there any objections if we release it ? (there were none). (JS) Note that reports for this meeting can be sent to the mailing list wlcg-scod@cernNOSPAMPLEASE.ch

Wednesday

EIS-on-Duty report:

SMOD report:

DBMOD report:

Monitoring / dashboard report:

Release update:

Questions from sites:

AOB:

Thursday

EIS-on-Duty report:

SMOD report:

DBMOD report:

Monitoring / dashboard report:

Release update:

Questions from sites:

AOB:

Friday

EIS-on-Duty report:

SMOD report:

DBMOD report:

Monitoring / dashboard report:

Release update:

Questions from sites:

AOB:

Edit | Attach | Watch | Print version | History: r2 < r1 | Backlinks | Raw View | WYSIWYG | More topic actions
Topic revision: r2 - 2008-01-16 - HarryRenshall
 
    • Cern Search Icon Cern Search
    • TWiki Search Icon TWiki Search
    • Google Search Icon Google Search

    LCG All webs login

This site is powered by the TWiki collaboration platform Powered by PerlCopyright & 2008-2019 by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding TWiki? Send feedback